# Matérn Kernels for Tunable Implicit Surface Reconstruction

### v1

@@ Abstract
We propose to use the family of Matérn kernels for tunable implicit surface reconstruction, building upon the recent success of kernel methods for 3D reconstruction of oriented point clouds. As we show, both, from a theoretical and practical perspective, Matérn kernels have some appealing properties which make them particularly well suited for surface reconstruction—outperforming state-of-the-art methods based on the arc-cosine kernel while being significantly easier to implement, faster to compute, and scaleable. Being stationary, we demonstrate that the Matérn kernels’ spectrum can be tuned in the same fashion as Fourier feature mappings help coordinate-based MLPs to overcome spectral bias. Moreover, we theoretically analyze Matérn kernel’s connection to Siren networks as well as its relation to previously employed arc-cosine kernels. Finally, based on recently introduced Neural Kernel Fields, we present data-dependent Matérn kernels and conclude that especially the Laplace kernel (being part of the Matérn family) is extremely competitive, performing almost on par with state-of-the-art methods in the noise-free case while having a more than five times shorter training time. [^1] Code available at: https://github.com/mweiherer/matern-surface-reconstruction.
@@ 1 Introduction
Recovering the shape of objects from sparse or only partial observations is a challenging task. Formally, let $\Omega\subset\mathbb{R}^{d}$ and $\mathcal{X}=\{x_{1},x_{2},\dots,x_{m}\}\subset\Omega$ be a set of $m$ points in $d$ dimensions which forms, together with associated per-point normals, a data set $\mathcal{D}=\mathcal{X}\times\{n_{1},n_{2},\dots,n_{m}\}\subset\Omega\times\mathbb{R}^{d}$. The goal of surface reconstruction is to recover the objects’ shape from $\mathcal{D}$. The shape of objects may be represented as dense point clouds, polygonal meshes, manifold atlases, voxel grids, or (as the zero-level set of) implicit functions, which is the representation we chose to focus on in this work. Specifically, if $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ denotes an implicit function, such as a signed distance function (SDF), then, from a practical perspective, implicit surface reconstruction aims at finding an optimal solution $\hat{f}$ to the following kernel ridge regression (KRR) problem:
@@
where the function space $H:=H(\Omega)$ is usually taken to be a reproducing kernel Hilbert space (RKHS) with associated reproducing kernel $k:\Omega^{2}\longrightarrow\mathbb{R}$, and $\lambda>0$. Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987). It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq. (1) turns into the following constrained minimization problem
@@
which makes it easier to see how $\hat{f}$ (thus the estimated surface $\mathcal{\hat{S}}$), behaves away from the input points, $\mathcal{X}$. Clearly, while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing. The properties of $\|f\|_{H}$ in $H$ are uniquely determined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq. (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$. Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$. This shows that the behavior of the norm in Eq. (2) and hence, of estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by choosing different kernel functions, effectively allowing us to inject certain inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
@@
Although already used in the mid-90s and early 2000s (Savchenko et al., 1995; Carr et al., 1997; 2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) eventually evolving as the new state-of-the-art. While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (Williams et al., 2021; 2022) employ the first-order arc-cosine kernel. Introduced by Cho & Saul (2009), arc-cosine kernels have been shown to mimic the computation of (two-layer) fully-connected ReLU networks; indeed, if the layer-width tends to infinity, the first-order arc-cosine kernel is identified with the Neural Tangent Kernel (NTK; Jacot et al. (2018)) of the network.
@@
Figure 2: Tunable Matérn kernels lead to better surface reconstructions compared to the previously employed arc-cosine kernel. Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
One important aspect of this kernel is that it does not have adjustable parameters. While this can be advantageous in some situations, it also prevents users from tuning their reconstructions. Clearly, if the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
@@
In this work, we propose to use a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2. Contrary to arc-cosine kernels, Matérn kernels are stationary (hence rotation- and translation-invariant), locally-supported (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace (for $\nu=1/2$) and Gaussian kernel (as $\nu\rightarrow\infty$). As we will show, both, from a theoretical and practical perspective, Matérn kernels have some intriguing properties which make them particularly well suited for surface reconstruction. Most notably, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details). In addition, our theoretical analysis reveals that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activation function; Sitzmann et al. (2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs. Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (i.e., the Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the hypersphere $\mathbb{S}^{d-1}$.
@@
Our experimental evaluation demonstrates that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable. Moreover, by leveraging Neural Kernel Fields (Williams et al., 2022), we show that data-dependent (a.k.a. learnable) Matérn kernels perform almost on par with highly sophisticated and well-engineered NKSR in the noise-free case while being more than five times faster to train.
@@ 2 Related Work
Early approaches for kernel-based 3D surface reconstruction mostly employ radial basis functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005). Nowadays, the probably most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; (Kazhdan & Hoppe, 2013)) which, however, can itself be viewed as kernel method as shown in Williams et al. (2021). Only recently, Neural Splines (NS; Williams et al. (2021)) proposed to use the so-called (first-order) arc-cosine kernel,
@@
for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques as well as non-linear methods based on neural networks by a large margin. This method laid the cornerstone for the Neural Kernel Field (NKF; Williams et al. (2022)) framework which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016). Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn a set of spatially-varying basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense. Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose surface reconstruction method that is highly scaleable and robust against noise, eventually becoming state-of-the-art.
3 Background: Matérn Kernels
We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) instead of the arc-cosine kernel for implicit surface reconstruction, being defined as:
@@ 3 Background: Matérn Kernels
where $\nu>0$ is a smoothness parameter that explicitly controls the differentiability, and $h>0$ is known as the shape parameter (or bandwidth) of the kernel. $\Gamma$ denotes the Gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$. For $\nu=p+\frac{1}{2}$, $p\in\mathbb{N}^{+}$ (i.e., if $\nu$ is a half-integer), Eq. (4) simplifies to a product of an exponential and polynomial of degree $p$, that is
$\Phi_{p+\frac{1}{2}}(\tau)=\exp\left(-\frac{\sqrt{2p+1}\tau}{h}\right)\frac{p!}{(2p)!}\sum_{i=0}^{p}\frac{(p+i)!}{i!(p-i)!}\left(\frac{2\sqrt{2p+1}\tau}{h}\right)^{p-i}.$ (5)
Matérn kernels generalize a variety of other kernels; using Eq. (5), the most popular ones can be written in closed-form as:
$p$ $=0:\Phi_{\text{1/2}}(\tau)=\exp\left(-\frac{\tau}{h}\right),$ (6)
$p$ $=1:\Phi_{\text{3/2}}(\tau)=\exp\left(-\frac{\sqrt{3}\tau}{h}\right)\left(1+\frac{\sqrt{3}\tau}{h}\right),$ (7)
$p$ $=2:\Phi_{\text{5/2}}(\tau)=\exp\left(-\frac{\sqrt{5}\tau}{h}\right)\left(1+\frac{\sqrt{5}\tau}{h}+\frac{5\tau^{2}}{3h^{2}}\right),$ (8)
@@
The spectral density of Matérn kernels is given as the inverse Fourier transform of $\Phi_{\nu}$, that is
$p_{\nu}(\omega)=\mathcal{F}^{-1}[\Phi_{\nu}]=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$ (10)
which is due to the fact that Matérn kernels are stationary (or shift-invariant—they only depend on the difference $x-y$ between two points, $x,y\in\Omega$); hence, kernel and associated spectral density are Fourier duals (a consequence of Bochner’s Theorem). Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$.
The RKHS induced by Matérn kernels is quite well understood (see, e.g., (Kanagawa et al., 2018)) and can be formulated as $H_{\nu}=\{f\in L_{2}(\mathbb{R}^{d})\cap C(\mathbb{R}^{d}):\|f\|_{H_{\nu}}^{2}<\infty\}$, where the RKHS norm is
$\|f\|_{H_{\nu}}^{2}=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega.$ (11)
For more information about Matérn kernels, we refer the interested reader to Porcu et al. (2023) for a recent review. Next, we revisit some basic properties of Matérn kernels.
@@ 3.1 Basic Properties
Differentiable. Matérn kernels allow for controlled smoothness, being exactly $\lceil\nu\rceil-1$ times differentiable (in the mean-square sense); for instance, while the Matérn 1/2 kernel is zero times differentiable, the Gaussian kernel (for $\nu\rightarrow\infty$) is infinitely often differentiable.
Stationary. Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$. If the distance is Euclidean, Matérn kernels are also isotropic (or homogeneous); they only depend on $\|x-y\|$. As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of $\mathcal{X}$. We note that, while the previously used arc-cosine kernel is also rotation invariant, it is not translation invariant.
Locally supported. As opposed to arc-cosine kernels, Matérn kernels are locally supported (or compact); that is, they tend to zero as the distance $\|x-y\|$ becomes large. This is a very desirable property, especially from a computational perspective, as it leads to sparse Gram matrices when multiplied with the kernel $\Phi^{\prime}_{\nu^{\prime}}(\tau)=\max\left\{0,\left(1-\tau/h^{\prime}\right)^{\nu^{\prime}}\right\}$ for suitable $\nu^{\prime}$ and $h^{\prime}$, effectively allowing the use of highly efficient and scalable sparse linear solvers. Please see Genton (2001) on how to choose suitable values for $\nu^{\prime}$ and the cut-off distance, $h^{\prime}$.
@@
Figure 3: Eigenvalue decay of Matérn kernels. While we fix $h=1$ and vary $\nu$ on the left, the EDR for $\nu=1/2$ and different values of $h$ is shown on the right. Larger values of $\nu$ and $h$, i.e., smoother kernels, lead to faster eigenvalue decay; hence slow convergence to high-frequency details.
4 Matérn Kernels for Tunable Surface Reconstruction

### v2

@@ Abstract
We propose to use the family of Matérn kernels for implicit surface reconstruction, building upon the recent success of kernel methods for 3D reconstruction of oriented point clouds. As we show from a theoretical and practical perspective, Matérn kernels have some appealing properties which make them particularly well suited for surface reconstruction—outperforming state-of-the-art methods based on the arc-cosine kernel while being significantly easier to implement, faster to compute, and scalable. Being stationary, we demonstrate that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias. Moreover, we theoretically analyze Matérn kernels’ connection to Siren networks as well as their relation to previously employed arc-cosine kernels. Finally, based on recently introduced Neural Kernel Fields, we present data-dependent Matérn kernels and conclude that especially the Laplace kernel (being part of the Matérn family) is extremely competitive, performing almost on par with state-of-the-art methods in the noise-free case while having a more than five times shorter training time. [^1] Code available at: https://github.com/mweiherer/matern-surface-reconstruction.
@@ 1 Introduction
Recovering the shape of objects from sparse or only partial observations is a challenging task. Formally, let $\Omega\subset\mathbb{R}^{d}$ and $\mathcal{X}=\{x_{1},x_{2},\dots,x_{m}\}\subset\Omega$ be a set of $m$ points in $d$ dimensions which forms, together with associated per-point normals, a dataset $\mathcal{D}=\mathcal{X}\times\{n_{1},n_{2},\dots,n_{m}\}\subset\Omega\times\mathbb{R}^{d}$. The goal of surface reconstruction is to recover the object’s shape from $\mathcal{D}$. The shape of objects may be represented as dense point clouds, polygonal meshes, manifold atlases, voxel grids, or (as the zero-level set of) implicit functions, which is the representation we chose to focus on in this work. Specifically, if $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ denotes an implicit function, such as a Signed Distance Function (SDF), then, from a practical perspective, implicit surface reconstruction aims at finding an optimal solution $\hat{f}$ to the following Kernel Ridge Regression (KRR) problem:
@@
where the function space $H:=H(\Omega)$ is usually taken to be a Reproducing Kernel Hilbert Space (RKHS) with associated reproducing kernel $k:\Omega\times\Omega\longrightarrow\mathbb{R}$, and $\lambda>0$. Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987). It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq. (1) turns into the constrained minimization problem,
@@
which makes it easier to see how the predicted surface behaves away from the input points. As seen from Eq. (2), while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing. The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq. (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$. Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$. This shows that the behavior of the norm and hence, estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by the kernel function, enabling easy injection of inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
@@
Although already used in the mid-90s and early 2000s (Savchenko et al., 1995; Carr et al., 1997; 2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) eventually evolving as the new state-of-the-art. While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (Williams et al., 2021; 2022) employ the first-order arc-cosine kernel. Introduced by Cho & Saul (2009), arc-cosine kernels have been shown to mimic the computation of (two-layer) fully-connected ReLU networks; indeed, if the layer-width tends to infinity, the first-order arc-cosine kernel is identified with the Neural Tangent Kernel (NTK; Jacot et al. (2018)) of the network. One important aspect of this kernel is that it does not have adjustable parameters. While this can be advantageous in some situations, it also prevents users from tuning their reconstructions. If the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
@@
Figure 2: Tunable Matérn kernels lead to better surface reconstructions than the previously employed arc-cosine kernel. Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
@@
In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2. Contrary to arc-cosine kernels, Matérn kernels are stationary (hence translation invariant), spatially decaying (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace and Gaussian kernel. As we will show, Matérn kernels have appealing properties, making them the ideal candidate for kernel-based surface reconstruction. From a practical perspective, we demonstrate that a simple change of the kernel function from arc-cosine to Matérn leads to a consistently improved reconstruction accuracy and a significant speedup. From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details). In addition, we prove that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; Sitzmann et al. (2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs. Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the sphere. In summary, our key contributions are:
  • We propose to use Matérn kernels for tunable implicit surface reconstruction.
  • We theoretically analyze Matérn kernels, relating them to Fourier feature mappings, Siren networks, and arc-cosine kernels. Moreover, we derive practical insights into how to choose the tunable bandwidth parameter based on a new bound of the $L_{2}$ reconstruction error.
  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (Williams et al., 2022).
@@
Our experimental evaluation reveals that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable. In addition to geometry, we show that Matérn kernels surpass the arc-cosine kernel in reconstructing other high-frequency scene attributes, such as texture. Finally, we demonstrate that learnable Matérn kernels (1) outperform the data-dependent arc-cosine kernel (as implemented in the original NKF framework) while being more than four times faster to train, and (2) perform almost on par with highly sophisticated and well-engineered NKSR in the noise-free case while having a more than five times shorter training time.
@@ 2 Related Work
Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005). Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; Kazhdan & Hoppe (2013)) which, however, can itself be viewed as kernel method (Williams et al., 2021). Only recently, Neural Splines (NS; Williams et al. (2021)) proposed to use the so-called (first-order) arc-cosine kernel,
@@
for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques and non-linear methods based on neural networks by a large margin. This method laid the cornerstone for Neural Kernel Fields (NKFs; Williams et al. (2022)) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016). Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense. Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
3 Matérn Kernels for Tunable Surface Reconstruction
We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) for implicit surface reconstruction, being defined as
@@ 3 Matérn Kernels for Tunable Surface Reconstruction
where $\nu>0$ is a smoothness parameter that explicitly controls the differentiability, and $h>0$ is known as the shape parameter (or bandwidth) of the kernel. $\Gamma$ denotes the Gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$. Matérn kernels generalize a variety of other kernel functions; the most popular ones can be written in closed form as
$\nu$ $=1/2:\Phi_{\text{1/2}}(\tau)=\exp\left(-\frac{\tau}{h}\right),$ (5)
$\nu$ $=3/2:\Phi_{\text{3/2}}(\tau)=\exp\left(-\frac{\sqrt{3}\tau}{h}\right)\left(1+\frac{\sqrt{3}\tau}{h}\right),$ (6)
$\nu$ $=5/2:\Phi_{\text{5/2}}(\tau)=\exp\left(-\frac{\sqrt{5}\tau}{h}\right)\left(1+\frac{\sqrt{5}\tau}{h}+\frac{5\tau^{2}}{3h^{2}}\right),$ (7)
@@
For more information about Matérn kernels, please refer to Porcu et al. (2023). Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
@@ 3.1 Basic Properties
Differentiable. Matérn kernels allow for controlled smoothness, being exactly $\lceil\nu\rceil-1$ times differentiable (in the mean-square sense). Since functions $f$ in an RKHS $H$ inherit the differentiability class of the inducing kernel $k$ due to the reproducing property, the from Eq. (1) reconstructed surface $\mathcal{\hat{S}}$ enjoys the same smoothness properties as $k$. In the context of 3D surface reconstruction, this allows an easy injection of inductive biases into the reconstruction problem; for instance, if the roughness or noisiness of the objects to be reconstructed is known in advance, one can adjust the smoothness of the kernel accordingly. This is not possible with the arc-cosine kernel.
Stationary. Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$. If the distance is Euclidean, Matérn kernels are also isotropic; they only depend on $\|x-y\|$. As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of points $\mathcal{X}$ (based on the reproducing property, translating input points does not change the reconstructed surface). Consequently, since the kernel’s value only depends on the relative positions, objects in a scene (or parts of an object) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same. This is in contrast to the arc-cosine kernel, whose value depends on the absolute position of points, see Eq. (3); hence, it is not translation invariant (thus non-stationary), and multiple instances of an object in a scene will generally not yield the same surface reconstruction. See Appendix A for further discussion.
Spatially decaying. As opposed to arc-cosine kernels, Matérn kernels are spatially decaying; they tend to zero as the distance $\|x-y\|$ becomes large. Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (Porcu et al., 2023), they can be considered effectively compact. From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (Genton, 2001), allowing the use of highly efficient and scalable sparse linear solvers. Finally, we note that simply truncating any kernel does not yield a valid, positive definite kernel in general.
@@
Figure 3: Eigenvalue decay of Matérn kernels. While we fix $h=1$ and vary $\nu$ on the left, the EDR for $\nu=1/2$ and different values of $h$ is shown on the right. Larger values of $\nu$ and $h$, i.e., smoother kernels, lead to faster eigenvalue decay; hence, slow convergence to high-frequency details.